# The electromagnetic field in Poisson gauge theory: the groupoidal approach

### v1

@@ Abstract
We consider the problem of defining the field strength of abelian potentials when the spacetime is a Poisson manifold, within the groupoidal approach. The natural definition in terms of gauge invariant momenta is proved to be equivalent to covariant and invariant tensors of a local symplectic groupoid representing a symplectic realization of the Poisson manifold. A Poisson Chern-Simons model is then proposed and its equations of motion are shortly discussed.
@@
2 Instituto de Ciencias Matemáticas ICMAT (CSIC-UAM-UC3M-UCM), Campus de Cantoblanco UAM, Calle Nicolás Cabrera, 13-15, 28049 Madrid, España.
@@ 3 Centro de Matemática, Computação e Cognição Universidade Federal do ABC, Santo André, SP, Brasil
6fcosmo[at]math.uc3m.es, 7vladislav.kupriyanov[at]gmail.com, 8patrizia.vitale[at]unina.it
@@ 1 Introduction
In recent years a novel approach to understand the behavior of gauge symmetries in the presence of non-commutativity has been introduced by two of the authors [1]. The idea is to first clarify at the semiclassical level, namely in the presence of a non-trivial Poisson structure, what compatibility requirements are necessary to have a proper action of the gauge algebra on the space of fields. Two main requirements were adopted: the commutative limit of the deformed gauge theory should reproduce the standard gauge theory and the algebra of infinitesimal gauge transformations should close. Within this framework several results have been obtained for the $U(1)$ symmetry [3, 2, 4, 5, 7, 6, 8, 9]. The derived models, dubbed Poisson gauge models, are field theories defined on a background which is a Poisson manifold, with a non-trivial Poisson bracket among space-time variables, depending on a parameter, say $\theta$. The quantization of these brackets corresponds to a fully noncommutative spacetime, whereas the limit of zero Poisson bracket ($\theta\rightarrow 0$) yields back the classical spacetime, hence the adjective "semiclassical" for this kind of models.
@@ 2 The groupoidal approach to Poisson gauge theory
The aim of this section is to give a brief introduction to symplectic groupoids, finalised to the application to Poisson gauge theory. We also seek to provide explicit expressions in local coordinates.
After recalling the main mathematical structures involved, we will start with two specific examples: the symplectic groupoid $\mathbb{R}^{n}\times\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ that integrates the canonical Poisson structure $\{x^{i},x^{j}\}=\Theta^{ij}$ on the base manifold $X=\mathbb{R}^{n}$ and the symplectic groupoid $\mathcal{G}\rightrightarrows X$ with $\mathcal{G}$ being the cotangent bundle $T^{\ast}G=G\times\mathfrak{g}^{\ast}$ integrating the Poisson manifold $X=\mathfrak{g}^{\ast}$, the dual Lie algebra of the Lie group $G$, with linear bracket $\{x^{i},x^{j}\}=c^{ij}_{k}x^{k}$. Then, we will consider the generalization to local symplectic groupoids, which have a close relationship to symplectic realizations.
@@ 2.1 Canonical non-commutativity, $\{x^{i},x^{j}\}=\Theta^{ij}=const$
According to def. 2.1, the elements of a groupoid can be composed, but the multiplication is only defined for those elements belonging to the set
$\mathcal{G}^{(2)}=\left\{\left((x^{\prime},p^{\prime}),(x,p)\right)\in\mathcal{G}\times\mathcal{G}\mid\mathsf{s}(x^{\prime},p^{\prime})=\mathsf{t}(x,p)\right\}$
of composable pairs. Then, the multiplication of two composable elements $\sigma_{1}=(x^{i},p_{i})$ and $\sigma_{2}=(x^{i}+\Theta^{ij}p_{j},p^{\prime}_{i})$ is given by,
@@
endowing $\mathcal{G}$ with a non-degenerate Poisson structure $\omega^{-1}$. This is a symplectic realization of the Poisson structure (2.24) and we recognize $\mathcal{G}\equiv\mathfrak{g}^{*}\times G$. The inverse of $\omega^{-1}$ defines a closed two-form $\omega$ which is a symplectic form on $\mathcal{G}$,

### v2

@@ Abstract
We consider the problem of defining the field strength of abelian potentials in Poisson Electrodynamics. Considering a local symplectic groupoid as a symplectic realization of the Poisson spacetime, it is possible to define several field strengths with different transformation properties. The introduction of gauge invariant momenta allows to find the relations that exist among all of them: in particular, they all vanish simultaneously. A Poisson Chern-Simons model is then proposed and its equations of motion are shortly discussed.
@@ 2 Centro de Matemática, Computação e Cognição Universidade Federal do ABC, Santo André, SP, Brasil
5fabio.di[at]uah.es, 6vladislav.kupriyanov[at]gmail.com, 7patrizia.vitale[at]unina.it
@@ 1 Introduction
In recent years a novel approach to understand the behavior of gauge symmetries in the presence of non-commutativity has been introduced by two of the authors [1]. The idea is to first clarify at the semiclassical level, namely in the presence of a non-trivial Poisson structure, what compatibility requirements are necessary to have a proper action of the gauge algebra on the space of fields. Two main requirements were adopted: the commutative limit of the deformed gauge theory should reproduce the standard gauge theory and the algebra of infinitesimal gauge transformations should close. Within this framework several results have been obtained for the $U(1)$ symmetry [3, 2, 4, 5, 7, 8, 6, 9, 10, 11]. The derived models, dubbed Poisson gauge models, are field theories defined on a background which is a Poisson manifold, with a non-trivial Poisson bracket among space-time variables, depending on a parameter, say $\theta$. The quantization of these brackets corresponds to a fully noncommutative spacetime, whereas the limit of zero Poisson bracket ($\theta\rightarrow 0$) yields back the classical spacetime, hence the adjective "semiclassical" for this kind of models.
@@ 2 The groupoidal approach to Poisson gauge theory
Symplectic groupoids were introduced at the end of the 80’s independently by Weinstein[13], Karasev[14] and Zkrezewski[15]. Their geometrical properties have been widely studied especially in connection with the problem of quantization: it has played an important role in the developments of geometric quantization of Poisson manifolds (see, for instance, [20, 21, 22]), in deformation quantization (see for instance [23, 24, 25, 26, 27]) or to provide geometric models of noncommutative algebras (see, for instance [16, 17, 18, 19]).
In this section we are not willing to provide a detailed review of this long history. Our goal, here, is to give a brief introduction to symplectic groupoids, aimed at the application to Poisson gauge theory. We also seek to provide explicit expressions in local coordinates which will be repeatedly used in the rest of the paper.
After recalling the main mathematical structures involved, we will focus on two specific examples: the symplectic groupoid $\mathbb{R}^{n}\times\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ that integrates the canonical Poisson structure $\{x^{i},x^{j}\}=\Theta^{ij}$ on the base manifold $X=\mathbb{R}^{n}$ and the symplectic groupoid $\mathcal{G}\rightrightarrows X$ with $\mathcal{G}$ being the cotangent bundle $T^{\ast}G=G\times\mathfrak{g}^{\ast}$ integrating the Poisson manifold $X=\mathfrak{g}^{\ast}$, the dual Lie algebra of the Lie group $G$, with linear bracket $\{x^{i},x^{j}\}=c^{ij}_{k}x^{k}$. Then, we will consider the generalization to local symplectic groupoids, which have a close relationship to symplectic realizations.
@@ 2.1 Canonical non-commutativity, $\{x^{i},x^{j}\}=\Theta^{ij}=const$
According to def. 2.1, the elements of a groupoid can be composed, but the multiplication is only defined for those elements belonging to the set $\mathcal{G}^{(2)}=\left\{\left((x^{\prime},p^{\prime}),(x,p)\right)\in\mathcal{G}\times\mathcal{G}\mid\mathsf{s}(x^{\prime},p^{\prime})=\mathsf{t}(x,p)\right\}$ of composable pairs. Then, the multiplication of two composable elements $\sigma_{1}=(x^{i},p_{i})$ and $\sigma_{2}=(x^{i}+\Theta^{ij}p_{j},p^{\prime}_{i})$ is given by,
@@
endowing $\mathcal{G}$ with a non-degenerate Poisson structure $\omega^{-1}$. This is a symplectic realization of the Poisson structure (2.24) and we recognize $\mathcal{G}\equiv\mathfrak{g}^{*}\times G$, where $G$ is a corresponding Lie group representing the momentum space. Note that the curved momentum space is a characteristic feature of noncommutative theories and their semiclassical counterparts. Other models with curved momentum space were considered in different contexts in [29, 30, 31, 32, 33, 34]. The inverse of $\omega^{-1}$ defines a closed two-form $\omega$ which is a symplectic form on $\mathcal{G}$,